# Lense-Thirring Acoustic Black Holes : Shadows and Light

### v1

@@ 4 Geometrical analysis
$r_{H}=2M,\quad r_{ac_{-}}=M\left(\xi-\sqrt{\xi(\xi-4)}\right),\quad r_{ac_{+}}=M\left(\xi+\sqrt{\xi(\xi-4)}\right),$ (4.1)
@@ 5 Shadows and null geodesics
In this section, we analyse the effective potential behavior of the slowly rotating acoustic black hole described by the metric (3.24). With the help of null geodesics, we derive relevant regions in which the effective potential has a particular behavior. Moreover, we investigate the shadows in the vicinity of an acoustic black hole and illustrate its behavior.
@@ 5.1 General approach
Studying the null geodesics is necessary in order to inspect the evolution of the massless particles surrounding the black hole under consideration [83]. In order to do so, we use the following Hamilton-Jacobi equation
$\frac{\partial S}{\partial\tau}=-\frac{1}{2}g^{\mu\nu}\frac{\partial S}{\partial x^{\mu}}\frac{\partial S}{\partial x^{\nu}},$ (5.1)
@@
$S=\frac{1}{2}m_{0}^{2}\tau-Et+L\phi+S_{r}\left(r\right)+S_{\theta}\left(\theta\right).$ (5.2)
The mass $m_{0}$ is equal to zero for the class of massless particles. The particles energy and angular momentum are denoted by $E$ and $L$, respectively, and the functions $S_{r}\left(r\right)$ and $S_{\theta}\left(\theta\right)$ only depend on $r$ and $\theta$, respectively. The following equation is obtained using the slow rotation metric (3.24) by substituting the Jacobi action (5.2) into the Hamilton-Jacobi equation (5.1)
$0=$ $-F(r)^{-1}\left(\frac{\partial S}{\partial t}\right)^{2}+\frac{1}{r^{2}}\left(\frac{\partial S}{\partial\phi}\right)^{2}-2a\sin^{2}\theta\frac{\left(1-F(r)\right)}{r^{2}F(r)}\left(\frac{\partial S}{\partial t}\right)\left(\frac{\partial S}{\partial\phi}\right)$ (5.3)
$+F(r)\left(\frac{dS_{r}}{dr}\right)^{2}+\frac{1}{r^{2}\sin^{2}\theta}\left(\frac{dS_{\theta}}{d\theta}\right)^{2}+\mathcal{O}\left(a^{2}\right).$ (5.4)
@@
$r^{4}F(r)\left(\frac{dS_{r}}{dr}\right)^{2}$ $=\frac{E^{2}r^{4}}{F(r)}+2aELr^{2}\frac{\left(1-F(r)\right)}{F(r)}-r^{2}\left(L^{2}+\mathcal{K}\right),$ (5.5)
$\frac{1}{\sin^{2}\theta}\left(\frac{dS_{\theta}}{d\theta}\right)^{2}$ $=\mathcal{K}-2aEL\frac{\left(1-F(r)\right)}{F(r)}\cos^{2}\theta.$ (5.6)
@@
$r^{2}\frac{dt}{d\tau}$ $=\frac{Er^{2}}{F(r)}-aL\frac{\left(1-F(r)\right)}{F(r)}\sin^{2}\theta,$ (5.7)
$r^{2}\frac{dr}{d\tau}$ $=\sqrt{R(r)},$ (5.8)
$r^{2}\frac{d\theta}{d\tau}$ $=\sqrt{\Theta(\theta)},$ (5.9)
$r^{2}\frac{d\phi}{d\tau}$ $=L+aE\frac{\left(1-F(r)\right)}{F(r)}\sin^{2}\theta.$ (5.10)
@@
$R(r)$ $=E^{2}r^{4}+2aELr^{2}\left(1-F(r)\right)-r^{2}F(r)(L^{2}+\mathcal{K}),$ (5.11)
$\Theta(\theta)$ $=\mathcal{K}\csc^{2}\theta-2aEL\frac{\left(1-F(r)\right)}{F(r)}\cot^{2}\theta.$ (5.12)
@@
$V_{eff}(r)=-\left(\frac{dr}{d\tau}\right)^{2}=-E^{2}+\frac{2aEL}{r^{2}}\left(F(r)-1\right)+\frac{F(r)}{r^{2}}\left(L^{2}+\mathcal{K}\right).$ (5.13)
@@
$\zeta=\frac{L}{E},\quad\eta=\frac{\mathcal{K}}{E^{2}}.$ (5.14)
These two impact parameters are used to rewrite the function $R(r)$ given in (5.11) as
$R(r)=E^{2}\left(r^{4}+2a\zeta r^{2}\left(1-F(r)\right)-r^{2}\,F(r)(\zeta^{2}+\eta)\right).$ (5.15)
@@
$R(r)\big|_{r_{i}}=0,\quad\frac{dR(r)}{dr}\big|_{r_{i}}=0.$ (5.16)
Thus, using (5.15) and (5.16), we obtain the impact parameters $\eta$ and $\zeta$, which are
$\zeta$ $=\frac{r\left(rF^{\prime}(r)-2F(r)\right)}{2aF^{\prime}(r)},$ (5.17)
$\eta$ $=\frac{r\left(8a^{2}F^{\prime}(r)-\left(rF^{\prime}(r)-2F(r)\right)\left(\left(r^{2}-4a^{2}\right)F^{\prime}(r)-2rF(r)\right)\right)}{4a^{2}F^{\prime 2}}.$ (5.18)
@@ 5.2 Photon and acoustic spheres
In this part of the paper, we use the equations established in the previous section to inspect the different regions that can be obtained from the effective potential of equation (5.13). To examine its behaviors and determine its maximum values, we illustrate such quantity versus $r$ for different values of the spin parameter $a$ and the acoustic parameter $\xi$ in figure (2).
@@
$V_{eff}^{\prime}(r)|_{r=r_{i}}=\frac{4aEL-\left(2F(r)-rF^{\prime}(r)\right)\left(2aEL+\mathcal{K}+L^{2}\right)}{r^{3}}=0.$ (5.20)
@@
$\frac{A\left(40M^{3}\xi-32M^{2}\xi r+6M(\xi+1)r^{2}-2r^{3}\right)+Br^{3}}{r^{6}}=\frac{-{\color[rgb]{0,0,1}e}r^{3}+br^{2}-cr+d}{r^{6}}=0,$ (5.21)
where ${\color[rgb]{0,0,1}e}=2A-B,b=6AM(\xi+1),c=32AM^{2}\xi,d=+40AM^{3}$. It can be shown that such an equation admits three solutions that are
$r_{1}$ $=\frac{b}{3e}+\frac{2^{1/3}\left(3ec-b^{2}\right)}{3e\left(X\right)^{1/3}}-\frac{\left(X\right)^{1/3}}{3\sqrt[3]{2}e},$ (5.22)
$r_{2}$ $=\frac{b}{3e}-\frac{\left(1+i\sqrt{3}\right)\left(3ec-b^{2}\right)}{3\ 2^{2/3}e\left(X\right)^{1/3}}+\frac{\left(1-i\sqrt{3}\right)\left(X\right)^{1/3}}{6\,2^{1/3}\,e},$ (5.23)
$r_{3}$ $=\frac{b}{3e}-\frac{\left(1-i\sqrt{3}\right)\left(3ec-b^{2}\right)}{3\ 2^{2/3}e\left(X\right)^{1/3}}+\frac{\left(1+i\sqrt{3}\right)\left(X\right)^{1/3}}{6\,2^{1/3}\,e},$ (5.24)
@@
$X=\sqrt{\left(27e^{2}d-9ebc+2b^{3}\right)^{2}-4\left(b^{2}-3ec\right)^{3}}-27e^{2}d+9ebc-2b^{3}.$ (5.25)
@@
$V_{eff}$ $=-\frac{4aELM+L^{2}(2M-r)+2KM+E^{2}r^{3}-Kr}{r^{3}},$ (5.26)
$V_{eff}^{\prime}$ $=\frac{12aELM+L^{2}(6M-2r)+6KM-2Kr}{r^{4}}.$ (5.27)
@@ 5.3 Shadow forms
As we have previously mentioned, in the frame of a LTABH, the optical shadow surrounding the event horizon, which describes the visual boundary that light cannot escape from by viewers, and the acoustic shadow, which describes the audible boundary of the sound waves detected by static listeners, may coexist. In what follows, we derive the equation governing such aspects of the black hole where the photon-phonon interaction has been neglected to avoid complications. Using equations (5.17), (5.18) and considering that the observer is located at the equatorial plan $\left(\theta=\frac{\pi}{2}\right)$, we illustrate the shadow radius denoted by $R_{s}$ as a function of the acoustic parameter $\xi$ in figure (4) for the different radii $r_{i}$ derived from the previous section, with $i=1,2,3$.
@@
$x$ $=\lim_{r_{*}\to\infty}\left(-r^{2}_{*}\sin^{2}\theta_{0}\frac{d\phi}{dr}\right),$ (5.28)
$y$ $=\lim_{r_{*}\to\infty}r^{2}_{*}\frac{d\theta}{dr},$ (5.29)
where $r_{*}$ is the distance between the black hole and the observer and $\theta_{0}$ is associated to the inclination angle between the line rotational axis of the black hole and the observer line of sight [84]. As a function of the impact parameters, these two celestial coordinates can be written as
$x$ $=-\zeta\csc\theta_{0},$ (5.30)
$y$ $=\sqrt{\eta-\zeta^{2}\cot^{2}\theta_{0}},$ (5.31)
@@ 6 Lense-Thirring effect and precession
In this part of the paper, we study the frame-dragging effect (i.e., Lense-Thirring precession) of the LTABH on the spin of a test gyroscope. One of the key predictions of Einstein’s general relativity is the Lense-Thirring precession, which has historically been the subject of enormous research efforts [85, 86]. The spin precession of a test gyroscope in the gravitational field of a spinning spherical body was first reported by Josef Lense and Hans Thirring in 1918. The Lense-Thirring precession effect can be used to anticipate the causal structures of the spacetime geometries. Consequently, it is crucial to understand the Lense-Thirring precession phenomenon in order to differentiate between the different black holes. In the following section, we review the mathematical description of the Lense-Thirring effect [87]. Then, we use the findings to analyse the frame-dragging in the vicinity of a LTABH.
@@ 6.1 Mathematical Framework
In the what follows, we examine a straightforward situation in which the gyroscopes are fixed at each point in space, and we look into how the precession frequency of the test gyroscope’s spin varies as the spatial points move. The gyroscope moves along an integral curve $\gamma(t)$ of a timelike Killing vector field $K$ since it is at rest in a stationary spacetime [88]. Then, the gyroscope’s four-velocity $u^{\alpha}$ can be expressed in terms of $K^{\alpha}$ as follows
$u^{\alpha}=\frac{1}{\sqrt{-K^{\beta}K_{\beta}}}K^{\alpha}.$ (6.1)
@@
$u^{\alpha}F_{\alpha}X^{\beta}=u^{\alpha}\nabla_{\alpha}X^{\beta}-X^{\alpha}a_{\alpha}u^{\beta}+X^{\alpha}u_{\alpha}a^{\beta},$ (6.2)
@@
$u^{\beta}\nabla_{\beta}e_{i}^{\alpha}=e_{i}^{\beta}a_{\beta}u^{\alpha}+\omega_{i}^{j}e_{j}^{\alpha},$ (6.3)
with $\omega_{i}^{j}=u^{\alpha}\nabla_{\alpha}e^{j}_{\beta}e^{\beta}_{i}$. With the help of the equations (6.2) and (6.3), the Fermi derivative component of the orthonormal basis in the direction of four-velocity $u_{\alpha}$ can be expressed as
$u^{\beta}F_{\beta}e_{\alpha}=\omega_{\alpha}^{\beta}e_{\beta}.$ (6.4)
@@
$\frac{d\mathcal{S}^{\alpha}}{d\tau}=\omega_{\beta}^{\alpha}\mathcal{S}^{\beta},$ (6.5)
@@
$\omega_{ij}=\epsilon_{ijk}\omega^{k},$ (6.6)
@@
$\omega_{ij}=\frac{1}{\sqrt{-K^{\alpha}K_{\alpha}}}e^{\beta}_{j}K^{\delta}\nabla_{\delta}e_{i\beta}.$ (6.7)
@@
$\Omega_{\mu}=\frac{1}{2\sqrt{-K^{\alpha}K_{\alpha}}}\ast\left(\overline{K}\wedge d\overline{K}\right)_{\mu},$ (6.8)
where $\wedge$ is the external product and $\ast$ is the Hodge operator [89]. The Lense-Thirring precession frequency for the spin vector $\mathcal{S}$ can be expressed as follows
$\Omega_{\mu}=\frac{1}{2\sqrt{-K^{\alpha}K_{\alpha}}}\mathcal{N}_{\mu}^{\nu\rho\sigma}K_{\nu}\partial_{\rho}K_{\sigma},$ (6.9)
@@
$ds^{2}=-g_{tt}dt^{2}+g_{rr}dr^{2}+g_{\theta\theta}d\theta^{2}+g_{\phi\phi}d\phi^{2}-2g_{t\phi}dtd\phi.$ (6.10)
@@
$\Omega=\left(g_{\theta\theta}\left(\Omega_{\theta}\right)^{2}+g_{rr}\left(\Omega_{r}\right)^{2}\right)^{\frac{1}{2}},$ (6.11)
where one has used the following equations [87]
$\Omega_{\theta}=\frac{g_{tt}}{2\sqrt{-g}}\left(\frac{g_{t\phi}}{g_{tt}}\right)_{,r},\quad\Omega_{r}=\frac{g_{tt}}{2\sqrt{-g}}\left(\frac{g_{t\phi}}{g_{tt}}\right)_{,\theta}.$ (6.12)

### v2

@@ 4 Geometrical analysis
$r_{H}=2M,\quad r_{ac_{-}}=M\left(\xi-\sqrt{\xi(\xi-4)}\right),\quad r_{ac_{+}}=M\left(\xi+\sqrt{\xi(\xi-4)}\right),$ (26)
@@ 5 Shadows and null geodesics
In this section, we analyse the effective potential behavior of the slowly rotating acoustic black hole described by the metric (24). With the help of null geodesics, we derive relevant regions in which the effective potential has a particular behavior. Moreover, we investigate the shadows in the vicinity of an acoustic black hole and illustrate its behavior.
@@ 5.1 General approach
Studying the null geodesics is necessary in order to inspect the evolution of the massless particles surrounding the black hole under consideration 1S1 . In order to do so, we use the following Hamilton-Jacobi equation
$\frac{\partial S}{\partial\tau}=-\frac{1}{2}g^{\mu\nu}\frac{\partial S}{\partial x^{\mu}}\frac{\partial S}{\partial x^{\nu}},$ (27)
@@
$S=\frac{1}{2}m_{0}^{2}\tau-Et+L\phi+S_{r}\left(r\right)+S_{\theta}\left(\theta\right).$ (28)
The mass $m_{0}$ is equal to zero for the class of massless particles. The particles energy and angular momentum are denoted by $E$ and $L$, respectively, and the functions $S_{r}\left(r\right)$ and $S_{\theta}\left(\theta\right)$ only depend on $r$ and $\theta$, respectively. The following equation is obtained using the slow rotation metric (24) by substituting the Jacobi action (28) into the Hamilton-Jacobi equation (27)
$0=$ $-F(r)^{-1}\left(\frac{\partial S}{\partial t}\right)^{2}+\frac{1}{r^{2}}\left(\frac{\partial S}{\partial\phi}\right)^{2}-2a\sin^{2}\theta\frac{\left(1-F(r)\right)}{r^{2}F(r)}\left(\frac{\partial S}{\partial t}\right)\left(\frac{\partial S}{\partial\phi}\right)$ (29)
$+F(r)\left(\frac{dS_{r}}{dr}\right)^{2}+\frac{1}{r^{2}\sin^{2}\theta}\left(\frac{dS_{\theta}}{d\theta}\right)^{2}+\mathcal{O}\left(a^{2}\right).$ (30)
@@
$r^{4}F(r)\left(\frac{dS_{r}}{dr}\right)^{2}$ $=\frac{E^{2}r^{4}}{F(r)}+2aELr^{2}\frac{\left(1-F(r)\right)}{F(r)}-r^{2}\left(L^{2}+\mathcal{K}\right),$ (31)
$\frac{1}{\sin^{2}\theta}\left(\frac{dS_{\theta}}{d\theta}\right)^{2}$ $=\mathcal{K}-2aEL\frac{\left(1-F(r)\right)}{F(r)}\cos^{2}\theta.$ (32)
@@
$r^{2}\frac{dt}{d\tau}$ $=\frac{Er^{2}}{F(r)}-aL\frac{\left(1-F(r)\right)}{F(r)}\sin^{2}\theta,$ (33)
$r^{2}\frac{dr}{d\tau}$ $=\sqrt{R(r)},$ (34)
$r^{2}\frac{d\theta}{d\tau}$ $=\sqrt{\Theta(\theta)},$ (35)
$r^{2}\frac{d\phi}{d\tau}$ $=L+aE\frac{\left(1-F(r)\right)}{F(r)}\sin^{2}\theta.$ (36)
@@
$R(r)$ $=E^{2}r^{4}+2aELr^{2}\left(1-F(r)\right)-r^{2}F(r)(L^{2}+\mathcal{K}),$ (37)
$\Theta(\theta)$ $=\mathcal{K}\csc^{2}\theta-2aEL\frac{\left(1-F(r)\right)}{F(r)}\cot^{2}\theta.$ (38)
@@
$V_{eff}(r)=-\left(\frac{dr}{d\tau}\right)^{2}=-E^{2}+\frac{2aEL}{r^{2}}\left(F(r)-1\right)+\frac{F(r)}{r^{2}}\left(L^{2}+\mathcal{K}\right).$ (39)
@@
$\zeta=\frac{L}{E},\quad\eta=\frac{\mathcal{K}}{E^{2}}.$ (40)
These two impact parameters are used to rewrite the function $R(r)$ given in (37) as
$R(r)=E^{2}\left(r^{4}+2a\zeta r^{2}\left(1-F(r)\right)-r^{2}\,F(r)(\zeta^{2}+\eta)\right).$ (41)
@@
$R(r)\big|_{r_{i}}=0,\quad\frac{dR(r)}{dr}\big|_{r_{i}}=0.$ (42)
Thus, using (41) and (42), we obtain the impact parameters $\eta$ and $\zeta$, which are
$\zeta$ $=\frac{r\left(rF^{\prime}(r)-2F(r)\right)}{2aF^{\prime}(r)},$ (43)
$\eta$ $=\frac{r\left(8a^{2}F^{\prime}(r)-\left(rF^{\prime}(r)-2F(r)\right)\left(\left(r^{2}-4a^{2}\right)F^{\prime}(r)-2rF(r)\right)\right)}{4a^{2}F^{\prime 2}}.$ (44)
@@ 5.2 Photon and acoustic spheres
In this part of the paper, we use the equations established in the previous section to inspect the different regions that can be obtained from the effective potential of equation (39). To examine its behaviors and determine its maximum values, we illustrate such quantity versus $r$ for different values of the spin parameter $a$ and the acoustic parameter $\xi$ in figure (2).
@@
$V_{eff}^{\prime}(r)|_{r=r_{i}}=\frac{4aEL-\left(2F(r)-rF^{\prime}(r)\right)\left(2aEL+\mathcal{K}+L^{2}\right)}{r^{3}}=0.$ (46)
@@
$\frac{A\left(40M^{3}\xi-32M^{2}\xi r+6M(\xi+1)r^{2}-2r^{3}\right)+Br^{3}}{r^{6}}=\frac{-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}e}r^{3}+br^{2}-cr+d}{r^{6}}=0,$ (47)
where ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}e}=2A-B,b=6AM(\xi+1),c=32AM^{2}\xi,d=+40AM^{3}$. It can be shown that such an equation admits three solutions that are
$r_{1}$ $=\frac{b}{3e}+\frac{2^{1/3}\left(3ec-b^{2}\right)}{3e\left(X\right)^{1/3}}-\frac{\left(X\right)^{1/3}}{3\sqrt[3]{2}e},$ (48)
$r_{2}$ $=\frac{b}{3e}-\frac{\left(1+i\sqrt{3}\right)\left(3ec-b^{2}\right)}{3\ 2^{2/3}e\left(X\right)^{1/3}}+\frac{\left(1-i\sqrt{3}\right)\left(X\right)^{1/3}}{6\,2^{1/3}\,e},$ (49)
$r_{3}$ $=\frac{b}{3e}-\frac{\left(1-i\sqrt{3}\right)\left(3ec-b^{2}\right)}{3\ 2^{2/3}e\left(X\right)^{1/3}}+\frac{\left(1+i\sqrt{3}\right)\left(X\right)^{1/3}}{6\,2^{1/3}\,e},$ (50)
@@
$X=\sqrt{\left(27e^{2}d-9ebc+2b^{3}\right)^{2}-4\left(b^{2}-3ec\right)^{3}}-27e^{2}d+9ebc-2b^{3}.$ (51)
@@
$V_{eff}$ $=-\frac{4aELM+L^{2}(2M-r)+2KM+E^{2}r^{3}-Kr}{r^{3}},$ (52)
$V_{eff}^{\prime}$ $=\frac{12aELM+L^{2}(6M-2r)+6KM-2Kr}{r^{4}}.$ (53)
@@ 5.3 Shadow forms
As we have previously mentioned, in the frame of a LTABH, the optical shadow surrounding the event horizon, which describes the visual boundary that light cannot escape from by viewers, and the acoustic shadow, which describes the audible boundary of the sound waves detected by static listeners, may coexist. In what follows, we derive the equation governing such aspects of the black hole where the photon-phonon interaction has been neglected to avoid complications. Using equations (43), (44) and considering that the observer is located at the equatorial plan $\left(\theta=\frac{\pi}{2}\right)$, we illustrate the shadow radius denoted by $R_{s}$ as a function of the acoustic parameter $\xi$ in figure (4) for the different radii $r_{i}$ derived from the previous section, with $i=1,2,3$.
@@
$x$ $=\lim_{r_{*}\to\infty}\left(-r^{2}_{*}\sin^{2}\theta_{0}\frac{d\phi}{dr}\right),$ (54)
$y$ $=\lim_{r_{*}\to\infty}r^{2}_{*}\frac{d\theta}{dr},$ (55)
where $r_{*}$ is the distance between the black hole and the observer and $\theta_{0}$ is associated to the inclination angle between the line rotational axis of the black hole and the observer line of sight 2S1 . As a function of the impact parameters, these two celestial coordinates can be written as
$x$ $=-\zeta\csc\theta_{0},$ (56)
$y$ $=\sqrt{\eta-\zeta^{2}\cot^{2}\theta_{0}},$ (57)
@@ 6 Lense-Thirring effect and precession
In this part of the paper, we study the frame-dragging effect (i.e., Lense-Thirring precession) of the LTABH on the spin of a test gyroscope. One of the key predictions of Einstein’s general relativity is the Lense-Thirring precession, which has historically been the subject of enormous research efforts body1 ; body2 . The spin precession of a test gyroscope in the gravitational field of a spinning spherical body was first reported by Josef Lense and Hans Thirring in 1918. The Lense-Thirring precession effect can be used to anticipate the causal structures of the spacetime geometries. Consequently, it is crucial to understand the Lense-Thirring precession phenomenon in order to differentiate between the different black holes. In the following section, we review the mathematical description of the Lense-Thirring effect bodyT . Then, we use the findings to analyse the frame-dragging in the vicinity of a LTABH.
@@ 6.1 Mathematical Framework
In the what follows, we examine a straightforward situation in which the gyroscopes are fixed at each point in space, and we look into how the precession frequency of the test gyroscope’s spin varies as the spatial points move. The gyroscope moves along an integral curve $\gamma(t)$ of a timelike Killing vector field $K$ since it is at rest in a stationary spacetime body3 . Then, the gyroscope’s four-velocity $u^{\alpha}$ can be expressed in terms of $K^{\alpha}$ as follows
$u^{\alpha}=\frac{1}{\sqrt{-K^{\beta}K_{\beta}}}K^{\alpha}.$ (60)
@@
$u^{\alpha}F_{\alpha}X^{\beta}=u^{\alpha}\nabla_{\alpha}X^{\beta}-X^{\alpha}a_{\alpha}u^{\beta}+X^{\alpha}u_{\alpha}a^{\beta},$ (61)
@@
$u^{\beta}\nabla_{\beta}e_{i}^{\alpha}=e_{i}^{\beta}a_{\beta}u^{\alpha}+\omega_{i}^{j}e_{j}^{\alpha},$ (62)
with $\omega_{i}^{j}=u^{\alpha}\nabla_{\alpha}e^{j}_{\beta}e^{\beta}_{i}$. With the help of the equations (61) and (62), the Fermi derivative component of the orthonormal basis in the direction of four-velocity $u_{\alpha}$ can be expressed as
$u^{\beta}F_{\beta}e_{\alpha}=\omega_{\alpha}^{\beta}e_{\beta}.$ (63)
@@
$\frac{d\mathcal{S}^{\alpha}}{d\tau}=\omega_{\beta}^{\alpha}\mathcal{S}^{\beta},$ (64)
@@
$\omega_{ij}=\epsilon_{ijk}\omega^{k},$ (65)
@@
$\omega_{ij}=\frac{1}{\sqrt{-K^{\alpha}K_{\alpha}}}e^{\beta}_{j}K^{\delta}\nabla_{\delta}e_{i\beta}.$ (66)
@@
$\Omega_{\mu}=\frac{1}{2\sqrt{-K^{\alpha}K_{\alpha}}}\ast\left(\overline{K}\wedge d\overline{K}\right)_{\mu},$ (67)
where $\wedge$ is the external product and $\ast$ is the Hodge operator body4 . The Lense-Thirring precession frequency for the spin vector $\mathcal{S}$ can be expressed as follows
$\Omega_{\mu}=\frac{1}{2\sqrt{-K^{\alpha}K_{\alpha}}}\mathcal{N}_{\mu}^{\nu\rho\sigma}K_{\nu}\partial_{\rho}K_{\sigma},$ (68)
@@
$ds^{2}=-g_{tt}dt^{2}+g_{rr}dr^{2}+g_{\theta\theta}d\theta^{2}+g_{\phi\phi}d\phi^{2}-2g_{t\phi}dtd\phi.$ (69)
@@
$\Omega=\left(g_{\theta\theta}\left(\Omega_{\theta}\right)^{2}+g_{rr}\left(\Omega_{r}\right)^{2}\right)^{\frac{1}{2}},$ (70)
where one has used the following equations bodyT
$\Omega_{\theta}=\frac{g_{tt}}{2\sqrt{-g}}\left(\frac{g_{t\phi}}{g_{tt}}\right)_{,r},\quad\Omega_{r}=\frac{g_{tt}}{2\sqrt{-g}}\left(\frac{g_{t\phi}}{g_{tt}}\right)_{,\theta}.$ (71)